OpenStudy (31356):

-3(y-5)=15 solve for y MEDAL REWARDED!!

OpenStudy (anonymous):

y=0

OpenStudy (anonymous):

y=0 -3y +15=15 -3y +15 -15= 15-15 -3y= 0

OpenStudy (31356):

Thx!

OpenStudy (31356):

Guruslimm because he showed the work. I wish I can give you a medal too froggyfroggy.

OpenStudy (anonymous):

it is okay people never give me the medals i am use to it

OpenStudy (31356):

Sorry